OpenStudy (anonymous):

Find the product. (n^3)^2 · (n^5)^4 is it n^14?

OpenStudy (anonymous):

No..

OpenStudy (saifoo.khan):

Now try.

OpenStudy (anonymous):

\(\large n^6.n^{20} = n^{26}\)

OpenStudy (usukidoll):

try to understand the methods of doing this type of problem. Order of Operations. Parenthesis exponents multiplication division addition subtraction n^3)^2 · (n^5)^4 (n^3)^2 means write (n^3) TWICE (n^3)(n^3) = N^[3+3] which is n^6 (n^5)^4 means write (n^5) FOUR times (n^5)(n^5)(n^5)(n^5) = n^5+5+5+5 = n^20 add them together WHOOOPS forgot that + and - doesn't exist when doing exponents so it's like this (n^6)(n^20) = n^ 6+20 = n^26